Williams Inc. produces a single product, a part used in the manufacture of automobile transmissions. Known for its quality and performance, the part is sold to luxury auto manufacturers around the world. Because this is a quality product, Williams has some flexibility in pricing the part. The firm calculates the price using a variety of pricing methods and then chooses the final price based on that information and other strategic information. A summary of the key cost information follows. Williams expects to manufacture and sell 58,500 parts in the coming year. While the demand for Williams’s part has been growing in the past 2 years, management is not only aware of the cyclical nature of the automobile industry, but also concerned about market share and profits during the industry’s current downturn. Total Costs Variable manufacturing $ 4,663,000 Variable selling and administrative 838,650 Facility-level fixed overhead 2,328,875 Fixed selling and administrative 658,495 Batch-level fixed overhead 343,000 Total investment in product line 22,333,000 Expected sales (units) 58,500 Required:

1. Determine the price for the part using a markup of 39% of full manufacturing cost.

2. Determine the price for the part using a markup of 20% of full life-cycle cost.

3. Determine the price for the part using a desired gross margin percentage to sales of 41%.

4. Determine the price for the part using a desired life-cycle cost margin percentage to sales of 28%.

5. Determine the price for the part using a desired before-tax return on investment of 13%.

6. Determine the total contribution margin and total operating profit for each of the methods in requirements 1 through 5.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Williams Inc.
Answer 1 Amount $
Variable manufacturing costs      4,663,000.00
Facility-level fixed overhead      2,328,875.00
Batch-level fixed overhead         343,000.00
Full manufacturing cost      7,334,875.00
Markup at 39%      2,860,601.25
Total sales 10,195,476.25
Units produced            58,500.00
Sell price                 174.28
Answer 2 Amount $
Variable manufacturing costs      4,663,000.00
Facility-level fixed overhead      2,328,875.00
Batch-level fixed overhead         343,000.00
Variable selling and administrative         838,650.00
Fixed selling and administrative         658,495.00
Full cycle cost      8,832,020.00
Markup at 20%      1,766,404.00
Total sales 10,598,424.00
Units produced            58,500.00
Sell price                 181.17
Answer 3
Gross margin percentage to sales 41%
It means if sales is 100
Then Gross margin is 41
And cost is (100-41)= 59
Markup on cost= 41/59= 69.492%
Full manufacturing cost      7,334,875.00
Markup at 69.492%      5,097,116.53
Total sales 12,431,991.53
Units produced            58,500.00
Sell price                 212.51
Answer 4
Life-cycle cost margin to sales 28%
It means if sales is 100
Then Life-cycle cost margin is 28
And Life-cycle cost is (100-28)= 72
Markup on Life-cycle cost cost= 28/72= 38.889%
Full cycle cost      8,832,020.00
Markup at 69.492%      3,434,674.44
Total sales 12,266,694.44
Units produced            58,500.00
Sell price                 209.69
Answer 5
Total investment    22,333,000.00
Before-tax return on investment 13%
Before-tax return on investment      2,903,290.00
Add: Full cycle cost      8,832,020.00
Total sales 11,735,310.00
Units produced            58,500.00
Sell price                 200.60
Answer 6 Answer 1 Answer 2 Answer 3 Answer 4 Answer 5
Total sales    10,195,476.25 10,598,424.00    12,431,991.53 12,266,694.44 11,735,310.00
Less: Variable costs
Variable manufacturing costs      4,663,000.00     4,663,000.00      4,663,000.00     4,663,000.00     4,663,000.00
Variable selling and administrative         838,650.00        838,650.00         838,650.00        838,650.00        838,650.00
Total Variable costs      5,501,650.00     5,501,650.00      5,501,650.00     5,501,650.00     5,501,650.00
Total contribution margin      4,693,826.25     5,096,774.00      6,930,341.53     6,765,044.44     6,233,660.00
Less: Fixed costs
Facility-level fixed overhead      2,328,875.00     2,328,875.00      2,328,875.00     2,328,875.00     2,328,875.00
Batch-level fixed overhead         343,000.00        343,000.00         343,000.00        343,000.00        343,000.00
Fixed selling and administrative         658,495.00        658,495.00         658,495.00        658,495.00        658,495.00
Total Fixed costs      3,330,370.00     3,330,370.00      3,330,370.00     3,330,370.00     3,330,370.00
Operating profit      1,363,456.25     1,766,404.00      3,599,971.53     3,434,674.44     2,903,290.00